# CASA: Class-Agnostic Shared Attributes in Vision-Language Models for Efficient Incremental Object Detection

### v1

@@ 4.2 Implementation Details
[…]
4.3 Results and Analyses
Table 2: CASA results (%) on COCO 2017 in two-phase setting 40+40. The best performance is highlighted in bold.
@@ 4.3 Results and Analyses
Table 3: CASA results ($AP$/$AP_{.5}$, %) on COCO 2017 in multi-phase setting. The tasks are categorized into two distinct scenarios: 40+10+10+10+10 and 40+20+20. Our method utilizes OWL-ViT as the baseline, while others are based on deformable DETR. The best performance is highlighted in bold.
@@
4.3.1 Two-phase setting.
In the Two-phase setting, we randomly divide the 80 classes of the COCO dataset into two experimental settings: 70+10 and 40+40. In Tab 1 and Tab 2 we compare the performance of our method, CASA, with other incremental object detection methods in terms of the metrics $AP$, $AP_{.5}$ and $AP_{.75}$ in both experimental settings. Additionally, in the 70+10 setting, we compare the difference in $FPP$ between our method and the previously best incremental object detection method, demonstrating that our method not only prevents forgetting but also achieves better performance than the previous task. As for False Positives ($FP$), in the 70+10 setting our method CASA has 31052 errors, which demonstrate an clear advantage, reducing at least 5000 errors than other methods. This indicates that our method effectively overcomes the issue of background shift compared to other incremental object detection methods. Based on the OWL-ViT model, in the first phase, the performance on 70 classes of object detection is slightly better than the previous optimal baseline Deformable DETR, while it is slightly inferior to Deformable DETR in the 40 classes setting. However, both in 70+10 and 40+40 settings, CASA consistently outperforms the state-of-the-art, better than CIOD and other incremental object detection approaches. In the 70+10 setting, our method achieves a 1.5% improvement in $AP_{.5}$ and a 1.9% improvement in $FPP$ over the current best methods, which is significantly more evident than the performance gains in the 40+40 setting. This is because our method CASA leverages class-agnostic shared attribute information for incremental object detection. The more classes learned in the first phase, the richer the preserved shared attribute information, which benefits the subsequent phases of incremental learning.
Table 4: Ablation experiments results of different components on COCO 2017 in multi-phase setting 40+10+10+10, with the best results highlighted in bold. The “—” indicates that the model lacks object detection capability from the first task, making it meaningless to calculate the FPP value in this context.
@@
4.3.2 Multi-phase setting.
In the multiple phases setting, we conducted experiments with configurations of 40+20+20 and 40+10+10+10+10 respectively. As before, the categories for each phase are randomly assigned from the COCO dataset’s 80 categories. Tab 3 records the $AP$ and $AP_{.5}$ at each step. We observe that our method, CASA, exhibits the best performance among incremental object detection methods in both multiple-phase experimental settings, even though the detection performance in the first phase is inferior to that of the method based on Deformable DETR. Notably, from the results of the 40+10+10+10 experiment, our method CASA demonstrates strong continuous learning capabilities across multiple steps, with $AP$ and $AP_{.5}$ significantly improved by 4.7% and 5%, respectively, compared to the current state-of-the-art method. This effectively indicates that our method can be applied to real-world scenarios for continuous learning.
4.4 Class-agnostic Shared Attributes
Our method, CASA, achieves incremental object detection through class-agnostic shared attributes in vision-language model. Here, we illustrate the process of increasing shared attributes using the 70+10 experimental setting as an example. In the first phase, we assume that each category is represented by 25 attribute information. In the first task, CASA only select 1314 attributes out of 2895 possible attributes for 70 categories, indicating that many attributes are shared among different categories within the first phase. In the second phase, CASA learns 10 additional categories, and in practice only 155 new attributes are added. This is partly due to some attributes being shared among these 10 categories, and more importantly, the attributes of these 10 categories are already used among the previous 70 categories, achieving efficient attribute sharing between the previous and current tasks.
@@ 4.4 Class-agnostic Shared Attributes
4.5 Ablation Experiments
We conduct ablation studies from two perspectives. First, we perform ablation experiments on our three main modules: updating the assignment matrix $A^{t}$ for attribute selection, adapting attribute embedding, and refining attribute embedding. As shown in Table 5 and Table 4, results show that all three components are essential. The first and second modules are indispensable, as the absence of either module leads to the model’s inability to perform effective detection. The inclusion of the refining module enhances the model’s detection results. Specifically, although in certain cases, methods that do not involve refinement may result in a slightly lower $FPP$, the detection metrics for each task, such as $AP$, $AP_{.5}$ and $AP_{.75}$, will show significant discrepancies compared to methods that do include refinement, and these discrepancies will continue to widen over time. Therefore, refinement remains necessary.
Table 5: Ablation experiments results of different components on COCO 2017 in two-phase setting 70+10, with the best results highlighted in bold. The “—” indicates that the model lacks object detection capability from the first task, making it meaningless to calculate the FPP value in context.
@@ 4.5 Ablation Experiments
On the other hand, we conduct ablation experiments focused on incremental learning. Specifically, we experimented with not sharing attribute information across the three modules and not applying the loss function that enforces the consistency of attribute embedding between the current phase and the previous phase. The results indicate that sharing attribute information during training and applying the loss function that controls phase-to-phase consistency lead to the best performance in incremental object detection.
4.6 Effect of Hyperparameters
The hyperparameter $\lambda_{1}$ serves as a regularization term for the assignment matrix. In our experiments, we fixed $\lambda_{1}$ at 0.01. We conducted extensive experiments on the COCO dataset using the 70+10 setting to analyze the effect of varying two hyperparameters $\lambda_{2}$ and $\lambda_{3}$, which control the consistency of the attribute embedding $\hat{E}_{a}^{t}[\ :Q,:]$ and $\hat{E}_{a}^{t-1}$ respectively. The results, as shown in the figure 3, demonstrate the variation in $AP_{.5}$ values when $\lambda_{2}$ and $\lambda_{3}$ are set to 0.5, 1, 2, and 4. We observed that the change in $AP_{.5}$ is less than 0.02%, indicating minimal impact. Compared to other incremental object detection methods, our approach achieves state-of-the-art performance.
Figure 3: Impact of the hyperparameters $\lambda_{2}$ and $\lambda_{3}$
4.7 Visualizations
The visualization results of incremental object detection are shown in the figure 4. The figure presents the visualization results under the 40+40 experimental setup on the COCO dataset. The red boxes represent the object classes learned by the model in the previous phase, while the green boxes indicate the object classes learned by the model in the current phase. From the visualization results, it can be observed that our method, CASA, demonstrates excellent detection performance for both the current and previously learned classes, effectively achieving incremental learning and overcoming background shift.
@@ 4.7 Visualizations
Figure 4: Visualization results for 40+40 setting on COCO 2017. The red boxes show object classes learned in the previous phase, while the green boxes represent those learned in the current phase.
5 Conclusions
@@
In this paper, we propose a novel method for incremental object detection (IOD), i.e., CASA, which effectively addresses the challenges of background shift by leveraging vision-language foundation models. By constructing a Class-Agnostic Shared Attribute base, CASA captures and retains common semantic attributes across incremental classes, ensuring consistent generalization across tasks. Our method preserves the original parameters of the pre-trained OWL-ViT model while incorporating parameter-efficient fine-tuning, which only adds 0.7% of parameter storage. This minimal increase in storage significantly enhances the scalability and adaptability of IOD. Extensive experiments conducted on the COCO dataset demonstrate that CASA achieves state-of-the-art performance in both two-phase and multi-phase incremental learning scenarios.

### v3

@@ IV-B Implementation Details
IV-C Results and Analyses
TABLE II: CASA results (%) on COCO 2017 in two-phase setting 40+40. The best performance is highlighted in bold.
@@ IV-C Results and Analyses
TABLE III: CASA results ($AP$/$AP_{.5}$, %) on COCO 2017 in multi-phase setting. The best performance is highlighted in bold.
@@
TABLE IV: Ablation results in multi-phase setting. The “—” indicates lacking object detection capability in the first task.
@@
IV-C1 Two-phase setting
We randomly divide the 80 classes of the COCO dataset into two experimental settings: 70+10 and 40+40. In Tab I and Tab II we compare the performance of our method, CASA, with other IOD methods in terms of the metrics $AP$, $AP_{.5}$ and $AP_{.75}$ in both experimental settings. Additionally, in the 70+10 setting, we compare the difference in $FPP$ between our method and the previously best IOD method, demonstrating that our method not only prevents forgetting but also achieves better performance than the previous task. As for False Positives ($FP$), in the 70+10 setting CASA has 31052 errors, which demonstrate an clear advantage, reducing at least 5000 errors than other methods. This indicates that our method effectively overcomes the issue of background shift compared to other IOD methods. Both in 70+10 and 40+40 settings, CASA consistently outperforms the state-of-the-art, better than CIOD and other IOD methods. In the 70+10 setting, our method achieves a 1.5% improvement in $AP_{.5}$ and a 1.9% improvement in $FPP$ over the current best methods, which is significantly more evident than the performance gains in the 40+40 setting. This is because CASA leverages class-agnostic shared attribute information for incremental object detection. The more classes learned in the first phase, the richer the preserved shared attribute information, which benefits the subsequent phases of incremental learning.
IV-C2 Multi-phase setting
We conduct experiments in 40+20+20 and 40+10+10+10+10 settings respectively. As before, the categories for each phase are randomly assigned. Tab III records the $AP$ and $AP_{.5}$ at each step. We observe that CASA exhibits the best performance among IOD methods in both multiple-phase settings, though the performance in the first phase is inferior to that of the method based on Deformable DETR. Notably, in the 40+10+10+10 setting, CASA demonstrates strong continuous learning capabilities across multiple steps, with $AP$ and $AP_{.5}$ significantly improved by 4.7% and 5%, respectively, compared to the current state-of-the-art method. This effectively indicates that our method can be applied to real-world scenarios for continuous learning.
IV-D Class-agnostic Shared Attributes
@@
Besides, The third part of supplementary materials IV provides an example of the process of increasing shared attributes.
IV-E Ablation Experiments
We perform ablation experiments on three modules: attributes selection and the assignment matrix $A^{t}$ updating, attributes adaption and attributes refinement. Table V and Table IV show that all three components are essential. Detailed explanation is in supplementary materials IV.
TABLE V: Ablation results in two-phase setting 70+10.
@@ IV-E Ablation Experiments
Specifically, we experiment with not sharing attribute across the three modules and not applying the loss function that enforces the consistency of attribute embedding between phases. Results indicate that sharing attribute information and applying the loss function that controls phase-to-phase consistency lead to the best performance in incremental object detection.
IV-F Effect of Hyperparameters and Visualizations
Figure 3 demonstrates the variation in $AP_{.5}$ values under the 70+10 setting when $\lambda_{1}$ and $\lambda_{2}$(described in supplemental materials III) are set to 0.5, 1, 2, and 4. We observed that the change in $AP_{.5}$ is less than 0.02%, indicating minimal impact($\lambda$ fixed at 0.01). The visualization results of IOD are shown in the figure 4. Detailed explanation is also in supplementary materials V.
@@ IV-F Effect of Hyperparameters and Visualizations
Figure 4: Visualization results for 40+40 setting. The red boxes show object classes learned in the previous phase, while the green boxes represent those learned in the current phase.
V Conclusions
@@
In this paper, we propose a novel method for IOD, CASA, which effectively addresses the challenges of background shift by leveraging vision-language foundation models. By constructing a Class-Agnostic Shared Attribute base, CASA captures and retains common semantic attributes across incremental classes. Our method preserves parameters of the pre-trained OWL-ViT model while incorporating parameter-efficient fine-tuning, which only adds 0.7% of parameter storage. Extensive experiments demonstrate that CASA achieves state-of-the-art performance in both two-phase and multi-phase incremental learning scenarios.
Supplemental Materials
I Details of attributes
I-A Shared Attributes Base
We use ten types of attributes, which are Color(such as red, yellow), Shape, Texture, Size, Context, Features, Appearance, Behavior, Environment, and Material. The attributeas base is generated only once. For each task, we selects a set of representive attributes using for detection from this attributes base according to their scores. For different tasks, the attributes used for detection are expanded, and shared among previous and new classes.
I-B Filtering out Unused Attributes
The initial Shared Attributes Base is generated with LLM. Some attributes are representative and effectively distinguish between different classes, but others, such as “shape is irregular,” do not significantly represent classes. We remove such attributes for two main reasons: on the one hand, this greatly reduces computational overhead; on the other hand, these attributes tend to have similar matching scores, which impact the subsequent process and reduce the accuracy of detection results.
When a new task arrives, we scores every attribute in the attribute base for new classes (including previously removed attributes). Attributes with high scores are retained, which may include previously selected attributes or previously removed ones. If an attribute is removed in a prior task but selected for a new task, our method appends these attributes to the end of the previous task’s remained attributes used for detection. Thus, the new set of attributes used for detection consists of attributes used in previous tasks and those newly selected in the current task.
I-C Process of Increasing Shared Attributes
Here, we illustrate the process of increasing shared attributes using the 70+10 experimental setting as an example. In the first phase, we assume that each category is represented by 25 attribute information. In the first task, CASA only select 1314 attributes out of 2895 possible attributes for 70 categories, indicating that many attributes are shared among different categories within the first phase. In the second phase, CASA learns 10 additional categories, and in practice only 155 new attributes are added. This is partly due to some attributes being shared among these 10 categories, and more importantly, the attributes of these 10 categories are already used among the previous 70 categories, achieving efficient attribute sharing between the previous and current tasks.
II Implementation Details Before Inference
II-A Attributes Adaption
After selecting the attribute embedding $\hat{E}_{a}^{t}$ based on $A^{t-1:t}$ for the current task, we also need to adapt the attribute embedding $\hat{E}_{a}^{t}$ to eliminate barriers between visual and textual information, because the textual attribute information generated from the text and the visual information are orthogonal in space, belonging to different domains. For each category in the current task, we take $M$ samples and calculate the visual mean embedding $\bar{E}^{v}$ for these $M$ samples:
$\bar{E}^{v}(c)=\frac{1}{M}\sum_{i=0}^{M}\mathbf{e}^{v}_{i}.$ (1)
$+\lambda_{1}\mathcal{L}_{MSE}\left(\hat{E}_{a}^{t}[\ :Q,:],\hat{E}_{a}^{t-1}\right),$
where $\lambda_{1}$ is a tunable hyperparameter. In this way, we adapt the attribute embedding $\hat{E}_{a}^{t}$, eliminating barriers between visual and textual information, which can be used for future refinement.
II-B Attributes Refinement
In Task $T_{t}$, after adapting the attribute information, we perform refining on a small scale to better apply the assignment matrix $A^{t}$ and the attribute embedding $\hat{E}_{a}^{t}$ to the subsequent inference. Similar to Equation 1 in the main text, we first compute the similarity vector $\hat{S}_{t}$ between the attribute embedding $\hat{E}_{a}^{t}$ and visual embedding $\mathbf{e}^{v}$ in the current task. Next, with the assignment matrix $A^{t-1:t}$, logits corresponding to these $C_{t}$ categories can be computed using Equation 2 in the main text. Unlike the previous process, at this stage, we need to keep the assignment matrix $A^{t}$ unchanged and update $\hat{E}_{a}^{t}$. We calculate the BCE loss between the probabilities $P$ of these categories and the targets $U$, which are a list of labels corresponding to the target categories in the image.
$\mathcal{L}_{ref}=$ $\ \mathcal{L}_{BCE}\left(P,U\right)+\lambda_{2}\mathcal{L}_{MSE}\left(\hat{E}_{a}^{t}[\ :Q,:],\hat{E}_{a}^{t-1}\right),$ (3)
where $\lambda_{2}$ is also a tunable hyperparameter. After refining the attribute embedding, the $\hat{E}_{a}^{t}$ and the $A^{t}$ can be used in the following inference stage.
III Hyperparameters Setting and parameter storage
III-A Hyperparameters Setting
[…]
III-B Parameter Storage
IV Ablation Experiments’ explanation
As shown in Table V and Table IV in the main text, results show that all three components(attributes selection and the assignment matrix $A^{t}$ updating, attributes adaption and attributes refinement) are essential. The first and second modules are indispensable, as the absence of either module leads to the model’s inability to perform effective detection. The inclusion of the refining module enhances the model’s detection results. Specifically, although in certain cases, methods that do not involve refinement may result in a slightly lower $FPP$, the detection metrics for each task, such as $AP$, $AP_{.5}$ and $AP_{.75}$, will show significant discrepancies compared to methods that do include refinement, and these discrepancies will continue to widen over time. Therefore, refinement remains necessary.
V Visualization details
Figure 4 in the main text presents the visualization results under the 40+40 experimental setup on the COCO dataset. The red boxes represent the object classes learned by the model in the previous phase, while the green boxes indicate the object classes learned by the model in the current phase. From the visualization results, it can be observed that our method, CASA, demonstrates excellent detection performance for both the current and previously learned classes, effectively achieving incremental learning and overcoming background shift.